Posada Company acquired 7,000 of the 10,000 outstanding shares of Sabathia Company on January 1, 2011, for $840,000. The subsidiary's total fair value was assessed at $1,200,000 although its book value on that date was $1,130,000. The $70,000 fair value in excess of Sabathia's book value was assigned to a patent with a 5-year remaining life. On January 1, 2013, Posada reported $1,085,000 equity method balance in the investment in Sabathia Company account. On October 1, 2013, Posada sells 1,000 shares of its investment for $191,000. During 2013, Sabathia reported net income of $120,000 and pad dividends of $40,000. These amounts are assumed to have occurred evenly throughout the year. a. What is the balance in investment account as of the October 2013 sale? b. Calculate the effect (profit or loss) from the sale of the 1,000 shares. c.) What is the balance in investment account after the sale ? d.) Calculate the non-controlling interest in Sabathia Co. as of December 31, 2015. Show calculations.
Answer #1

a. What is the balance in investment account as of the October 2013 before sale?

Answer : 1,140,650

Investment balance as on 1st Jan 2013        1,085,000
Add : Share of Profit for 9 month of 2013              63,000 =120000/12*9*70%
Less: Amortization of acquisition premium for 9 month                7,350 =70000/5*9/12*70%
Investment balance as on 1st Oct 2013        1,140,650
Less : Cost of Sales of Investment            162,950 =1140650/7000*1000
Balance after sale of Investment            977,700

b. Calculate the effect (profit or loss) from the sale of the 1,000 shares

Number of shares as on 1st Oct 2013          7,000.00
Value Per share (1,140,650/7,000)              162.95
Number of shares sold          1,000.00
Value of investment sold      162,950.00
Less : Sales consideration of investment sold      191,000.00
Gain on Sale of Investment        28,050.00

c.) What is the balance in investment account after the sale ?

Answer : 977,700

Investment balance as on 1st Jan 2013        1,085,000
Add : Share of Profit for 9 month of 2013              63,000 =120000/12*9*70%
Less: Amortisation of acquisition premium for 9 month                7,350 =70000/5*9/12*70%
Investment balance as on 1st Oct 2013        1,140,650
Less : Cost of Sales of Investment            162,950 =1140650/7000*1000
Balance after sale of Investment            977,700

d.) Calculate the non-controlling interest in Sabathia Co. as of December 31, 2015. Show calculations.

Answer : 646,400

Balance of Investment as on 1st Oct 2013 after sale of Investment            977,700
Add : Share of Profit for 3 month of 2013 18,000 =120000/12*3*60%
Less : Proportionate dividend received 24,000 =40,000*60%
Less: Amortization of acquisition premium for 3 month                2,100 =70000/5*3/12*60%
Balance of Investment as on 31st Dec 2013            969,600
Non-controlling interest share on 31st Dec 2013            646,400 993600/60*40